### v2

# Higher-Order Fermion Interactions in BCS Theory

ABSTRACT
We investigate the impact of higher-order fermionic deformations in multiflavor Bardeen-Cooper-Schrieffer (BCS) theory. Focusing specifically on the 8-fermion interaction, we show that this term can have significant consequences on the dynamics of the system. In certain regions of parameter space, the theory continues to exhibit second-order phase transitions with mean-field critical exponents and the same critical temperature; however, the temperature dependence of the superconducting gap can deviate markedly from conventional BCS behavior. In other regions, the theory exhibits first-order phase transitions. We conclude by discussing potential phenomenological applications of these theories.
1 Introduction
The Bardeen-Cooper-Schrieffer (BCS) theory of superconductivity remains as a major achievement of the 20th century Physics [1]. Through phonon interactions, electrons feel an effective quartic interaction which forces them to condense into Cooper pairs driving superconductivity. Yet, from the perspective of effective field theory, it remains surprising that a quartic fermionic interaction can produce such a pronounced effect. Naive power-counting suggests that the quartic coupling is an irrelevant interaction in spacetime dimensions greater than two. However, as argued in [2], the presence of the Fermi surface significantly modifies the naive scaling, making the quartic coupling marginally relevant in specific kinematic regimes. This observation naturally raises the question of how this effect might propagate in the presence of higher-order couplings. It is indeed reasonable to expect that the RG flow will mix the running of the various couplings, so that the infrared relevance of the quartic interaction will, in general, extend to and influence the behavior of the higher-order terms.
In this paper we set out to study this question in detail. For concreteness, we restrict our analysis to a class of models with $\psi^{4}$ and $\psi^{8}$ interactions. As we will see, depending on the coefficient of the higher-order fermionic coupling, the transition can be second order or first order. The transition is second order below a certain critical value of the $\psi^{8}$ coupling. In this case, the critical exponents still have mean-field values, but the gap curve is deformed as compared with the BCS prediction. In another range of parameter, the higher-order couplings drive the system into a regime where the phase transition becomes first-order. We will discuss potential phenomenological applications of the various scenarios in section 5.
2 The model
Our starting point is the familiar non-relativistic Lagrangian describing the BCS model, extended to $N$ fermionic flavors as in [3],
$\mathcal{L}=\sum_{a=1}^{N}\sum_{\sigma=\uparrow,\downarrow}\bar{\psi}_{a,\,\sigma}\Big{(}\partial_{\tau}-\frac{\vec{\partial}^{2}}{2m}-\mu\Big{)}\psi_{a,\,\sigma}-\frac{g}{N}\,(\sum_{a=1}^{N}\bar{\psi}_{a,\uparrow}\bar{\psi}_{a,\downarrow})\,(\sum_{b=1}^{N}\psi_{b,\downarrow}\psi_{b,\uparrow})\ .$ (2.1)
This model assumes a $SO(N)\times U(1)_{e}$ continuous global symmetry under which ${\psi}_{a,\uparrow}$ and $\bar{\psi}_{a,\downarrow}$ transform as $\mathbf{N}_{1}$ while ${\psi}_{a,\downarrow}$ and ${\bar{\psi}}_{a,\uparrow}$ transform as $\mathbf{\overline{N}}_{1}$. Under $U(1)_{e}$, ${\psi}_{a,\uparrow}$, ${\psi}_{a,\downarrow}$ have the same charge equal to 1-1, whereas $\bar{\psi}_{a,\uparrow}$, $\bar{\psi}_{a,\downarrow}$ have charges equal to 1.111There is also a $U(1)^{N}$ global symmetry rotating $(\psi_{a,\uparrow},\bar{\psi}_{a,\downarrow})\rightarrow(e^{i\theta_{a}}\,\psi_{a,\uparrow},e^{-i\theta_{a}}\,\bar{\psi}_{a,\downarrow})$ for each $a$. This symmetry will not play any role in the following.
Introducing a Hubbard-Stratonovich field $\Delta$ transforming in the $\mathbf{1}_{-2}$, one writes the equivalent Lagrangian
$\mathcal{L}^{\prime}=\sum_{a=1}^{N}\sum_{\sigma=\uparrow,\downarrow}\bar{\psi}_{a,\,\sigma}\Big{(}\partial_{\tau}-\frac{\vec{\partial}^{2}}{2m}-\mu\Big{)}\psi_{a,\,\sigma}-\Delta\,(\sum_{a=1}^{N}\bar{\psi}_{a,\uparrow}\bar{\psi}_{a,\downarrow})-\bar{\Delta}\,(\sum_{b=1}^{N}\psi_{b,\downarrow}\psi_{b,\uparrow})+\frac{N}{g}\Delta\bar{\Delta}\,.$ (2.2)
It is worth noting that, more generically, we could have considered an interaction $\Delta_{ab}\,\bar{\psi}_{a,\uparrow}\bar{\psi}_{b,\downarrow}+{\rm h.c.}$ exhibiting a $U(N)\times U(N)$ global symmetry; of which our model is a restriction. This generic class also includes the models describing the so-called Type 1.5 superconductors [4].
The equation of motion for $\Delta$ and $\bar{\Delta}$ sets
$\bar{\Delta}=\frac{g}{N}\,(\sum_{a=1}^{N}\bar{\psi}_{a,\uparrow}\bar{\psi}_{a,\downarrow})\,,\qquad\Delta=\frac{g}{N}\,(\sum_{b=1}^{N}\psi_{b,\downarrow}\psi_{b,\uparrow})\,.$ (2.3)
Substituting back into $\mathcal{L}^{\prime}$ we recover the original Lagrangian $\mathcal{L}$.
As described in the introduction, we are interested in exploring the effect of a priori irrelevant couplings. Thus, we now add a $\psi^{8}$ deformation. This term is non-vanishing for $N\geq 2$. The Lagrangian is
$\mathcal{L}=$ $\sum_{a=1}^{N}\sum_{\sigma=\uparrow,\downarrow}\bar{\psi}_{a,\,\sigma}\Big{(}\partial_{\tau}-\frac{\vec{\partial}^{2}}{2m}-\mu\Big{)}\psi_{a,\,\sigma}-\frac{g}{N}\,(\sum_{a=1}^{N}\bar{\psi}_{a,\uparrow}\bar{\psi}_{a,\downarrow})\,(\sum_{b=1}^{N}\psi_{b,\downarrow}\psi_{b,\uparrow})$
$-\frac{\lambda}{2N^{3}}\,(\sum_{a=1}^{N}\bar{\psi}_{a,\uparrow}\bar{\psi}_{a,\downarrow})^{2}(\sum_{b=1}^{2}\psi_{b,\downarrow}\psi_{b,\uparrow})^{2}\,.$ (2.4)
where $\lambda\geq 0$ for stability. Next, we introduce auxiliary fields $\Delta$ and $\Phi$ as follows
$\mathcal{L}^{\prime}=$ $\sum_{a=1}^{N}\sum_{\sigma=\uparrow,\downarrow}\bar{\psi}_{a,\,\sigma}\Big{(}\partial_{\tau}-\frac{\vec{\partial}^{2}}{2m}-\mu\Big{)}\psi_{a,\,\sigma}-\Delta\,(\sum_{a=1}^{N}\bar{\psi}_{a,\uparrow}\bar{\psi}_{a,\downarrow})-\bar{\Delta}\,(\sum_{b=1}^{N}\psi_{b,\downarrow}\psi_{b,\uparrow})+\frac{N}{g}\Delta\bar{\Delta}$
$-\lambda(\bar{\Phi}\Phi)\left(\bar{\Phi}\,(\sum_{a=1}^{N}\psi_{a,\downarrow}\psi_{a,\uparrow})+\Phi(\sum_{a=1}^{N}\bar{\psi}_{a,\uparrow}\bar{\psi}_{a,\downarrow})\right)+\frac{3}{2}\lambda N(\bar{\Phi}\Phi)^{2}\,.$ (2.5)
The equations of motion for $\bar{\Phi}$ and $\Phi$ set
$\Phi=\frac{1}{N}\,\left(\sum_{a=1}^{N}\psi_{a,\downarrow}\psi_{a,\uparrow}\right)\ ,\qquad\bar{\Phi}=\frac{1}{N}\,\left(\sum_{a=1}^{N}\bar{\psi}_{a,\uparrow}\bar{\psi}_{a,\downarrow}\right)\,.$ (2.6)
It should be stressed that section 2 is only one out of many possible models, as one could consider higher $\psi^{n}$. Also, for generic $N$, typically there exist various possible contractions of the flavor indices contributing to a given $\psi^{n}$ interaction. Due to the Grassman nature of the fermionic field, for sufficiently large $n$, the interaction terms vanish. For instance, for $N=1$ all $\psi^{n}$ interactions with $n>4$ vanish. For $N=2$ all $\psi^{n}$ terms with $n>8$ vanish and the $\psi^{8}$ interaction is unique, given by
$\frac{\lambda}{8}(\bar{\psi}_{1,\uparrow}\bar{\psi}_{1,\downarrow})\,(\psi_{1,\downarrow}\psi_{1,\uparrow})(\bar{\psi}_{2,\uparrow}\bar{\psi}_{2,\downarrow})\,(\psi_{2,\downarrow}\psi_{2,\uparrow})\ .$
For $N=2$, the Lagrangian (2) represents the most general Lagrangian consistent with the symmetries.
2.1 Integrating out the fermions
Writing the model as in section 2, the fermions appear quadratically and hence can be integrated out. Let us introduce
$\Psi_{a}=\left(\begin{array}[]{c}\psi_{a,\uparrow}\\ \bar{\psi}_{a,\downarrow}\end{array}\right)\,,\qquad\Psi^{\dagger}_{a}=(\bar{\psi}_{a,\uparrow},\,\psi_{a,\downarrow})\,.$ (2.7)
and
$\mathcal{A}=\left(\begin{array}[]{cc}-\partial_{\tau}+\frac{\vec{\partial}^{2}}{2m}+\mu&\Delta+\lambda(\bar{\Phi}\Phi)\Phi\\ \bar{\Delta}+\lambda(\bar{\Phi}\Phi)\bar{\Phi}&-\partial_{\tau}-\frac{\vec{\partial}^{2}}{2m}-\mu\end{array}\right)\,.$ (2.8)
Then section 2 becomes
$\mathcal{L}=\sum_{a=1}^{N}\Psi_{a}^{\dagger}\,\mathcal{A}\,\Psi_{a}+\frac{N}{g}\Delta\bar{\Delta}+\frac{3}{2}\lambda N(\bar{\Phi}\Phi)^{2}\,.$ (2.9)
Integrating out the $N$ fermions, the one-loop effective Lagrangian becomes
$\mathcal{L}_{\rm 1-loop}=N\,\left(-{\rm Tr}\log\mathcal{A}+\frac{1}{g}\Delta\bar{\Delta}+\frac{3}{2}\lambda(\bar{\Phi}\Phi)^{2}\right)\ .$ (2.10)
For the theory at finite temperature, ${\rm Tr}\log\mathcal{A}$ is given by
${\rm Tr}\log\mathcal{A}=T\,\sum_{n}\int\frac{d^{d}\vec{k}}{(2\pi)^{d}}\,\log\big{[}\big{(}(2n+1)\pi T\big{)}^{2}+\xi^{2}+|\Delta+\lambda(\bar{\Phi}\Phi)\Phi|^{2}\big{]}\,$, (2.11)
where
$\xi=\frac{\vec{k}^{2}}{2m}-\mu\,.$ (2.12)
Note that ${\rm Tr}\log\mathcal{A}=\mathcal{D}(X)$, with $X=|\Delta+\lambda(\bar{\Phi}\Phi)\Phi|^{2}$. The equations of motion for $\bar{\Delta}$, $\Delta$ are
$\Delta+\lambda(\bar{\Phi}\Phi)\Phi]\,\mathcal{D}^{\prime}(X)+\frac{\Delta}{g}=0\,\qquad-[\bar{\Delta}+\lambda(\bar{\Phi}\Phi)\bar{\Phi}]\,\mathcal{D}^{\prime}(X)+\frac{\bar{\Delta}}{g}=0\,.$ (2.13)
where we have introduced the notation $\mathcal{D}^{\prime}\equiv\partial\mathcal{D}/\partial X$. Let us now consider the equations for $\bar{\Phi},\ \Phi$. In view of (2.3), (2.6), one can anticipate that the solution will be $\bar{\Phi}=\bar{\Delta}/g$ , $\Phi=\Delta/g$ As usual, one can set $\bar{\Delta}=\Delta$ (the constant phase is irrelevant for the dynamics as it can be removed by a $U(1)$ rotation of the fermions). This sets $\bar{\Phi}=\Phi$. The remaining equation for $\Phi$ is
$\Phi^{2}[\Delta+\lambda\,\Phi^{3}]\,\mathcal{D}^{\prime}(X)+\Phi^{3}=0\,.$ (2.14)
There is a trivial solution $\Delta=\Phi=0$ representing the uncondensed phase. To look for the condensed phase, we assume $\Delta\neq 0$ and $\Phi\neq 0$. Then, combining (2.13) and (2.14), we find, as expected
$\Phi=\frac{1}{g}\,\Delta\,.$ (2.15)
We are finally then left with a single equation, which reads
$\left(-(1+c\,\Delta^{2})\,\mathcal{D}^{\prime}+\frac{1}{g}\right)\Delta=0\,.$ (2.16)
where
$c\equiv\frac{\lambda}{g^{3}}\ ,$ (2.17)
and $\mathcal{D}^{\prime}$ is evaluated at $\Phi$ given by (2.15).
The gap in the condensed phase is determined by the equation
$(1+c\,\Delta^{2})\,\mathcal{D}^{\prime}=\frac{1}{g}\ .$ (2.18)
In order to further proceed, we now need to compute $\mathcal{D}^{\prime}(X)$. Explicitly, it reads
$\mathcal{D}^{\prime}=T\,\sum_{n}\,\int\frac{d^{d}\vec{k}}{(2\pi)^{d}}\,\frac{1}{\big{(}(2n+1)\pi T\big{)}^{2}+\xi^{2}+\Delta^{2}\,\eta^{2}}\,$, (2.19)
where
$\eta\equiv 1+c\,\Delta^{2}\ .$ (2.20)
Performing the sum, one finds
$\mathcal{D}^{\prime}=\frac{1}{2}\,\int\frac{d^{d}\vec{k}}{(2\pi)^{d}}\,\frac{\tanh\frac{\sqrt{\xi^{2}+\Delta^{2}\,\eta^{2}}}{2T}}{\sqrt{\xi^{2}+\Delta^{2}\,\eta^{2}}}\,.$ (2.21)
Integrating, we find
$\mathcal{D}=\int\frac{d^{d}\vec{k}}{(2\pi)^{d}}\Big{[}\sqrt{\xi^{2}+\Delta^{2}\,\eta^{2}}+2T\,\log\frac{1}{2}\Big{(}1+e^{-\frac{\sqrt{\xi^{2}+\Delta^{2}\,\eta^{2}}}{T}}\Big{)}\Big{]}\,.$ (2.22)
This formula will be used in the calculation of the free energy.
In standard BCS theory, one assumes that the relevant dynamics takes place near the Fermi surface. Under this assumption, one can approximate
$\int\frac{d^{d}\vec{k}}{(2\pi)^{d}}\sim\nu\,\int_{-\omega_{*}}^{\omega_{\star}}d\xi\,$, (2.23)
where $\omega_{\star}$ is the Debye energy (playing the role of a physical cutoff) and $\nu$ is the density of states at the Fermi level, defined by a $\xi(k)=\frac{k^{2}}{2m}-\mu=0$. The final form of the gap equation is therefore:
$\frac{1}{g\,\nu\,(1+c\,\Delta^{2})}=\int_{0}^{\omega_{\star}}d\xi\,\frac{\tanh\frac{\sqrt{\xi^{2}+\Delta^{2}\,\eta^{2}}}{2T}}{\sqrt{\xi^{2}+\Delta^{2}\,\eta^{2}}}\,.$ (2.24)
The new terms with coefficients $b$ and $c$ represent corrections to standard BCS theory and cannot be absorbed into a redefinition of $\Delta$.
2.2 Free energy
The free energy is obtained by computing the action evaluated on the solution. This yields (we set $\Delta=\bar{\Delta}$)
$\frac{F}{N}=\frac{1}{g}\Delta^{2}+\frac{3}{2}\lambda\Phi^{4}-\int\frac{d^{d}\vec{k}}{(2\pi)^{d}}\Big{[}\sqrt{\xi^{2}+\Delta^{2}\,\eta^{2}}+2T\,\log\frac{1}{2}\Big{(}1+e^{-\frac{\sqrt{\xi^{2}+\Delta^{2}\,\eta^{2}}}{T}}\Big{)}\Big{]}\,.$ (2.25)
Choosing the uncondensed state –with free energy $F_{0}$– as the reference, the free-energy difference $f\equiv\frac{1}{N}(F-F_{0})$ is
$f=\frac{1}{g}\Delta^{2}+\frac{3\,c}{2\,g}\,\Delta^{4}-\int\frac{d^{d}\vec{k}}{(2\pi)^{d}}\Big{[}\big{(}\sqrt{\xi^{2}+\Delta^{2}\,\eta^{2}}-\xi\big{)}+2T\,\log\Big{(}\frac{1+e^{-\frac{\sqrt{\xi^{2}+\Delta^{2}\,\eta^{2}}}{T}}}{1+e^{-\frac{\xi}{T}}}\Big{)}\Big{]}\,$, (2.26)
Assuming the same approximation as in eq. 2.23 for the momentum integral we find
$f=\frac{1}{g}\Delta^{2}+\frac{3\,c}{2\,g}\,\Delta^{4}-2\nu\int_{0}^{\omega_{\star}}d\xi\Big{[}\big{(}\sqrt{\xi^{2}+\Delta^{2}\,\eta^{2}}-\xi\big{)}+2T\,\log\Big{(}\frac{1+e^{-\frac{\sqrt{\xi^{2}+\Delta^{2}\,\eta^{2}}}{T}}}{1+e^{-\frac{\xi}{T}}}\Big{)}\Big{]}\,.$ (2.27)
2.3 Fluctuations
To analyze the system we have so far used the saddle-point method applied to the one-loop action (2.10). Owing to the overall $N$ factor, this approximation becomes exact in the infinite $N$ limit. In order to examine the stability, we can study the fluctuations around the classical configuration. Writing the fields as $\Delta=\Delta_{0}+\delta\Delta$, $\Phi=\Phi_{0}+g^{-1}\delta\Phi$, where $\Delta_{0}$, $\Phi_{0}$ are the saddle-point values of the fields, one can systematically expand (2.10) in perturbation theory around the classical solution. To quadratic order, we can read off the mass matrix for the fluctuations. It is straightforward to check that such matrix has one zero eigenvalue, corresponding to the Goldstone mode arising from the spontaneous symmetry breaking of the $U(1)_{e}$. The rest of the eigenvalues are positive, corresponding to positive squared masses for the fluctuations. The explicit expressions are rather lengthy. For $c\ll 1$, to linear order in $c$, they simplify and read
$M^{2}=\{0,\,\frac{c\,\Delta_{0}^{2}}{g},\,-\mathcal{D}^{\prime\prime}\,\Delta_{0}^{2}+\frac{c\,\Delta_{0}^{2}}{2\,g}(1-4\,\mathcal{D}^{\prime\prime}\,g\,\Delta_{0}^{2}),\,\frac{3\,c\,\Delta_{0}^{2}}{2\,g}\}\,$, (2.28)
where
$\mathcal{D}^{\prime\prime}(X)=-\frac{\nu}{4}\int_{0}^{\omega_{\star}}\ d\xi\ \frac{\sinh\frac{\sqrt{\xi^{2}+X}}{T}-\frac{\sqrt{\xi^{2}+X}}{T}}{(\xi^{2}+X)^{\frac{3}{2}}\cosh^{2}\frac{\sqrt{\xi^{2}+X}}{2T}}\ .$ (2.29)
$\mathcal{D}^{\prime\prime}$ is given by minus the integral of a positive definite quantity. Therefore $\mathcal{D}^{\prime\prime}$ is negative, $\mathcal{D}^{\prime\prime}<0$, for any value of $X$. Thus, apart from the protected massless Goldstone mode, all fluctuations around the vacuum possess positive squared mass, ensuring the perturbative stability of the model.
3 Analysis of the condensed phase
In this section, an analytical study of the possible condensed phases is carried out by examining the gap equation eq. 2.24.
3.1 The $T=0$ limit
We first consider the small temperature region. As $T\rightarrow 0$, the $\tanh$ in eq. 2.24 goes to one, and the integral can be computed explicitly. One finds the following equation for $\Delta_{0}\equiv\Delta(T=0)$:
$\eta_{0}\,\Delta_{0}=\frac{\omega_{\star}}{\sinh(1/(g\,\nu\,\eta_{0}))}\ ,\qquad\eta_{0}\equiv 1+c\Delta_{0}^{2}\ .$ (3.1)
This is a transcendental equation for $\Delta$ that can be solved numerically. The BCS case $b=c=0$ gives an estimate, which also applies to case where $c$ is small,
$\Delta_{0}\approx\frac{\omega_{\star}}{\sinh(1/g\nu)}\ ,\qquad{\rm for}\ \ c\Delta_{0}^{2}\ll 1\ .$ (3.2)
3.2 Expansion near the critical temperature
Assuming that $\Delta(T)$ vanishes at a critical temperature $T_{c}$, we have
$\frac{1}{g\nu}=\int_{0}^{\frac{\omega_{\star}}{T_{c}}}\frac{\tanh\frac{x}{2}}{x}\approx\log\frac{k\omega_{\star}}{T_{c}}\,,\quad k\approx 1.134\qquad\leadsto\qquad T_{c}\approx k\,\omega_{\star}e^{-\frac{1}{g\nu}}\ .$ (3.3)
Note that the critical temperature $T_{c}$ at which the gap vanishes, $\Delta=0$, is the same as in BCS theory since in this case the correction with $c$ coefficient vanishes. However, because $\Delta$ can in principle be multivalued, the temperature $T_{c}$ may not be the critical temperature of the onset of the phase transition. In section 4 we will examine instances in which the system undergoes a first-order phase transition at a higher critical temperature $T_{c}^{\star}$.
The region close to $T_{c}$ is amenable to analytical treatment. Let us expand the gap equation around $T_{c}$. Defining $\delta T=T_{c}-T$ and neglecting $O(\Delta^{3})$ terms, we get
$\alpha_{1}\,\delta T-\left(\beta_{0}-\frac{c}{g\nu}\right)\,\Delta^{2}=0\,.$ (3.4)
where
$\alpha_{1}=\frac{1}{2T_{c}}\,\int_{0}^{\frac{\omega_{\star}}{T_{c}}}dx\,\frac{1}{\cosh^{2}\frac{x}{2}}\,,\qquad\beta_{0}=\frac{1}{4T_{c}^{2}}\,\int_{0}^{\frac{\omega_{\star}}{T_{c}}}dx\ \frac{1}{x^{3}\,\cosh^{2}\frac{x}{2}}\,\left(\sinh(x)-x\right)\,.$ (3.5)
It follows that
$\Delta^{2}=\frac{\alpha_{1}}{\beta_{0}-\frac{c}{g\nu}}\,\delta T\,.$ (3.6)
We recall that $c\geq 0$ for stability of the potential. Thus we have a second order transition provided $\beta_{0}-\frac{c}{g\nu}>0$, and first order otherwise. The critical $c$ happens at $c_{0}=g\nu\,\beta_{0}$. Explicitly
$c_{0}=\frac{1}{4T_{c}^{2}}\,\frac{\int_{0}^{\frac{\omega_{\star}}{T_{c}}}\frac{1}{x^{3}\,\cosh^{2}\frac{x}{2}}\,(\sinh x-x)}{\int_{0}^{\frac{\omega_{\star}}{T_{c}}}\frac{\tanh\frac{x}{2}}{x}}\approx n_{0}\,\frac{g\nu}{4T_{c}^{2}},\ \ n_{0}\approx 0.426\,$, (3.7)
where the integral in the numerator has been estimated assuming that $\omega_{\star}\gg T_{c}$. In order to study the behavior right at the critical coupling $c=c_{0}$, we need to keep the next order in the gap $\Delta(T)$, leading to
$\Delta^{4}=\gamma\,(T_{c}-T)+O(\delta T^{2})\ ,\qquad\gamma>0\ .$ (3.8)
Thus, at this critical coupling $c_{0}$, the critical exponent is $1/4$.
To provide further support for this picture, we can calculate the free energy near $T_{c}$. We find
$f\sim\frac{1}{2}\,\frac{\alpha_{1}^{2}\,\nu}{(\beta_{0}-\frac{c}{g\nu})}\,\delta T^{2}+\cdots\,.$ (3.9)
Thus we see that $f$ precisely changes sign at $c=c_{0}$ and goes from a negative to a positive concavity at $T_{c}$, consistent with the change from a second to a first order phase transition (cf. red, black and blue vs. green curves in fig. 2 below).
Just as in conventional BCS (type I) superconductors, the Meissner effect dictates the temperature dependence of the critical magnetic field. One has
$\frac{1}{2}H_{c}(T)^{2}=Nf(T)\ .$ (3.10)
For $b=0$ and $c<c_{0}$, near the critical temperature, using (3.9), one obtains the usual behavior
$H_{c}(T)\approx H_{c}(0)\left(1-\left(\frac{T}{T_{c}}\right)^{2}\right)\ .$ (3.11)
Away from the critical temperature, in regions where $\Delta$ is not small, there will be deviations from BCS, but the phase diagram will be qualitatively the same.
4 Numerical analysis of the condensed phase
In this section we numerically solve the gap equation and compute the free energy of the system for different values of the parameters.
According to the analysis of the previous section, the theory with $c<c_{0}$ should exhibit a second order phase transition with critical exponent $1/2$ ($1/4$ for the fine-tuned case $c=c_{0}$) and a first order phase transition in the regime $c>c_{0}$. This is supported by fig. 1, where we show the numerical solution to the gap equation for $\Delta$ as a function of the temperature for different values of $c$. For $c<c_{0}$ (red and black curves) the curve has a similar behavior to the standard BCS case. In turn, for $c>c_{0}$ (green curve), $\Delta$ develops a lump (in the case at hand, reaching up to $T_{0}\sim 4.5$) and becomes non-single valued in the range $T\in(T_{c},\,T_{0})$, indicating the presence of a first order phase transition in this regime.
Figure 1: $\Delta$ vs. $T$ for $c=0$ (standard BCS, in red), then $c=0.8\,c_{0}$ (black), $c=c_{0}$ (dashed, blue), $c=1.2\,c_{0}$ (green). Here $g\nu=0.3$, $\omega_{\star}=100$.
This scenario can be explicitly tested by computing the free energy in each case as shown in fig. 2. For $c>c_{0}$ the free energy exhibits the characteristic swallowtail of a first-order phase transition showing that at some $T_{c}^{\star}$ ($\sim 4.3$ in the example shown) there is a first order phase transition.
Refer to caption   Refer to caption
(a) (b)
Figure 2: a) Free energy vs. $T$ (same conventions as in figure 1). b) Enlarged view of the free energy in the case $c=1.2\,c_{0}$. The lower branch in Fig. 1 corresponds to the branch with positive free energy. The upper branch in figure 1 has negative free energy until $T^{\star}_{c}\approx 4.3$, where a first-order phase transition takes place. For $T>T^{\star}_{c}$, $\Delta=0$.
Although the critical exponents in the region $c<c_{0}$ remain the same as in mean-field theory and match those of the BCS theory, the temperature dependence of the gap exhibits a qualitatively distinct profile compared to the BCS case (cf. the black curve with the BCS curve in red in fig. 1).
5 Conclusions
In this paper we have studied the effect of $\psi^{4}$ and $\psi^{8}$ interactions interactions at finite temperature and finite chemical potential in the context of BCS-like phase transitions. While these interactions might naively be expected to play a minor role because these operators are “superficially” irrelevant operators, the present results indicate that they can significantly affect the system’s behavior, in some instances significantly modifying temperature dependence of the gap or leading to a first-order transition. The mechanism by which an irrelevant operator turns into a marginally relevant operator is well understood in BCS theory [2]. In the standard BCS theory, the effective quartic interaction is generated through electron-phonon interactions. Below a critical temperature, the quartic interaction forces fermions to condense into Cooper pairs driving superconductivity. Under naive dimensional analysis, the quartic coupling possesses a mass dimension of 2d2-d, rendering it irrelevant in spacetime dimensions $d>2$. Thus, from the standpoint of effective field theory, it may be unexpected that a quartic fermionic interaction can lead to condensation. Nevertheless, as shown in [2], the presence of a Fermi surface fundamentally modifies this scaling, rendering the quartic interaction marginally relevant within particular kinematic regimes. This observation naturally prompts a systematic investigation into the behavior of higher-order couplings and their potential role in mediating analogous effects. Although a comprehensive investigation lies beyond the scope of this work, it is reasonable to expect that, in the presence of a Fermi surface, the $\psi^{8}$ coupling will not have the naive running (dictated by dimensional analysis) under the Renormalization Group, since the loop integrals have a different behavior [2]. In addition, the RG flow intermixes and propagates the low energy behavior of the quartic fermion coupling to higher-order fermionic interactions.
When the $\psi^{8}$ interaction is turned off, the model describes BCS theory with $N$ fermion species and the usual $\psi^{4}$ interaction. In practical applications, models with several species of fermions arise in multiband superconductors [6], where fermions of different bands are treated as different species. Generically, higher fermion interactions may naturally appear in the effective Lagrangian for systems with $N\geq 2$. This suggests that these models might also be relevant in the description of multicomponent (in particular, type 1.5) superconductors (see [4] for a review).
Other potential applications of these models include superconducting systems with strong electron-phonon coupling. These can exhibit deviations from the BCS behavior. The gap curve is typically flatter for most of the temperature range and drops sharply near $T_{c}$ (see e.g. [7]). This is in accordance with the behavior exhibited by the models with $c<c_{0}$ (black curve in Fig. 1). Given this remarkable agreement, it would be highly interesting to investigate whether strong electron-phonon coupling underlies the effective $\psi^{8}$ interaction. A more fundamental derivation of the $\psi^{8}$ interaction could also provide valuable insight into estimating the coupling constant $c$ in these superconducting systems.
For $c>c_{0}$, the gap becomes double valued above $T_{c}$. This feature appears in a number of models, in particular, in the Eliashberg-Nambu theory for strong-coupling superconductors [8, 9, 10]. In this parameter range, the present models exhibit a first-order superconducting transition, where the gap edge drops from a finite value to zero. While most common superconductors have a continuous superconducting transition, there are cases where a discontinuous jump in the order parameter is expected. This includes multicomponent superconductors, strong-coupling Eliashberg superconductors and heavy fermions (first-order phase transitions also appear in the presence of magnetic fields).
Acknowledgements
JGR acknowledges financial support from the Spanish MCIN/AEI/10.13039/501100011033 grant PID2022-126224NB-C21. DRG is grateful to the University of Barcelona for hospitality. DRG is supported in part by the Spanish national grant MCIU-22- PID2021-123021NB-I00